Find the vector equation of the plane determined by the points (-1,-2,1),(1,0,1) and (1,-1,1). Also check whether (1/2,1/2,1/2) lies on it


1
Expert's answer
2021-05-24T02:01:02-0400

"\begin{vmatrix}
 x-(-1) & y-(-2) & z-1 \\
 1-(-1) & 0-(-2) & 1-1 \\
 1-(-1) & -1-(-2) & 1-1
\end{vmatrix}=0"


"\begin{vmatrix}
 x+1 & y+2 & z-1\\
 2 & 2 & 0\\
2 & 1 & 0
\end{vmatrix}=0"


"(z-1)[2-4]=0"

"z=1"


"\vec r.(0\vec i+0\vec j+1\vec k)=1"


Put the point (x = 1/2. y = 1/2, z = 1/2) on the equation of plane i.e

"0x+0y+z=1"

"\dfrac{1}{2}\cancel=1"

Therefore, the point (1/2,1/2,1/2) does not lie on the plane